Business Solutions sells upscale modular desk units and office chairs in the ratio of 3:2 (desk unit;chair). The selling price are $1,250 per desk unit and $500 per chair. The variable costs are $750 per desk unit and $250 per chair. Fixed costs are $120,000.

Required:

1. Compute the selling price per composite unit.

2. Compute the variable costs per compsite unit.

3. Compute the break-even point in composite units.

4. Compute the number of units of each product that would be sold at the break-even point.
Answer #1

1. Selling price per composite unit = 3*$1,250 + 2*$500

Selling price per composite unit = $3,750+1,000 = $4,750

2. Variable cost per composite unit = 3*$750 + 2*$250

Variable cost per composite unit = 2,250 + 500 = $2,750

3. Break even point in composite units = Fixed cost / Contribution per unit

Break even point in composite units = $120,000 / $4,750-2,750

Break even point in composite units = $120,000 / 2,000 = 60 units

4. Number of units of each product sold at break even point

Desk units = 60*3/5 = 36 units

Chairs = 60*2/5 = 24 chairs